OpenStudy (anonymous):

3x - y = -4 5x + 4y = -8

OpenStudy (anonymous):

What do we have to do ?

OpenStudy (anonymous):

find the points that it would be on the graph

OpenStudy (anonymous):

Umm it would be on a graph oh darn lol i thought a fraction

OpenStudy (anonymous):

Hero will be here to help you

OpenStudy (mertsj):

From the first equation y = 3x+4 Replace y in the second equation and solve. Can you do that this time?

OpenStudy (anonymous):

And some others that are a bit smart :P

OpenStudy (anonymous):

Solve simultaneously multiply Eq1 by4 and andd to Eq2 to get x

OpenStudy (anonymous):

x=-24/17

OpenStudy (mertsj):

\[5x+4(3x+4)=-8\]

OpenStudy (anonymous):

I have 17x + 16 = -8. What to do after that?

OpenStudy (mertsj):

Subtract 16 from both sides.

OpenStudy (anonymous):

for x I have -1.41

OpenStudy (mertsj):

that it correct but it might be better to leave it in it's exact form which is -24/17

OpenStudy (mertsj):

Because now you have to find y which is 3x+4

OpenStudy (mertsj):

\[y=3(-\frac{24}{17})+4\]

OpenStudy (mertsj):

\[y=\frac{-72}{17}+4\]

OpenStudy (mertsj):

\[y=\frac{-72}{17}+\frac{68}{17}=\frac{-4}{17}\]

OpenStudy (mertsj):

So the solution is (-24/17,-4/17)

OpenStudy (anonymous):

Thanks